OpenStudy (anonymous):

MEDAL FOR ALLL ANSWERSSSSSS Write the equation of the line in slope-intercept form that passes through the points (-2, 6) and (-1, 4).

OpenStudy (anonymous):

slope= (6-4)/(-2--1)=2 y=-2x+b 6=-2*-2+b 6=4+b 2=b y=-2*x+2

OpenStudy (anonymous):

sorry, that should be (6-4)/(-2--1)=-2

OpenStudy (anonymous):

mx+b

OpenStudy (anonymous):

haha i got a medal

OpenStudy (anonymous):

ok guys i give medals for all my answers in all my questions and thanks and can i get some medals lol

OpenStudy (anonymous):

question is closed no more medals

OpenStudy (anonymous):

answer random stuff on this http://openstudy.com/study#/updates/4f7b829ee4b0587805a10517 i will give u medal

OpenStudy (anonymous):

yay a metal lol

OpenStudy (anonymous):

i mean give me a metal for a good question lol

OpenStudy (anonymous):

im giving a free metal to people who can answer my question

OpenStudy (anonymous):

which fraction is equivalent to 1_3

OpenStudy (anonymous):

i y answer up 11 min ago lol and my friend is just anfused as me

OpenStudy (anonymous):

fast

OpenStudy (anonymous):

idc i just love to give out medals lol